# Hear-Your-Click: Interactive Video-to-Audio Generation via Object-aware Contrastive Audio-Visual Fine-tuning

### v1

@@ Abstract.
Video-to-audio (V2A) generation shows great potential in fields such as film production. Despite significant advances, current V2A methods, which rely on global video information, struggle with complex scenes and often fail to generate audio tailored to specific objects or regions in the videos. To address these limitations, we introduce Hear-Your-Click, an interactive V2A framework that enables users to generate sounds for specific objects in the videos by simply clicking on the frame. To achieve this, we propose Object-aware Contrastive Audio-Visual Fine-tuning (OCAV) with a Mask-guided Visual Encoder (MVE) to obtain object-level visual features aligned with corresponding audio segments. Furthermore, we tailor two data augmentation strategies: Random Video Stitching (RVS) and Mask-guided Loudness Modulation (MLM), aimed at enhancing the model’s sensitivity to the segmented objects. To effectively measure the audio-visual correspondence, we design a new evaluation metric, the CAV score, for evaluation. Extensive experiments demonstrate that our framework offers more precise control and improved generation performance across various metrics. Project Page: https://github.com/SynapGrid/Hear-Your-Click
@@
Figure 1. Overview of Hear-Your-Click.
[…]
After users select a single frame and interactively choose specific objects or areas, Hear-Your-Click generates the corresponding audio. This user-friendly approach empowers individuals to customize the audio according to their preferences and interests.
@@ 1. Introduction
Video-to-Audio (V2A) generation has shown great potential for applications in diverse domains, including film production, social media, and accessibility services. Recent research in the field has yielded significant advances, particularly with diffusion models demonstrating remarkable performance in cross-modal generation tasks (Ruiz et al., 2023; Huang et al., 2023; Evans et al., 2024; Ma et al., 2024; Ju et al., 2024).
@@
Despite considerable advances, several unresolved challenges persist in the field of Video-to-Audio (V2A) generation. One of the primary challenges is the comprehension of intricate semantic information within videos. Videos often contain multiple objects and their corresponding actions, requiring V2A models to generate appropriate sound effects for each object while accounting for their interactions. This complexity poses a significant challenge, as most V2A models struggle to handle such rich semantic information effectively.
@@
Another challenge is the lack of fine-grained control over the audio generation process. Existing methods predominantly rely on global video information, which is insufficient for customizing audio for specific objects or regions within the video according to user needs (Iashin and Rahtu, 2021; Luo et al., 2024; Xing et al., 2024). This limitation hinders the ability of the generated audio to meet the specific requirements of different scenarios, thereby restricting the practical applicability of V2A models in domains such as interactive media and film production. Introducing finer control mechanisms would not only enhance the accuracy of the generated audio but also improve the feasibility of V2A models in real-world applications.
@@
To address these challenges, this work proposes an interactive V2A generation framework named Hear-Your-Click. As shown in Fig. 1, by allowing users to select specific objects within a video with a single click, Hear-Your-Click generates sounds that correspond to the selected regions, ensuring precise synchronization and alignment with the visual content. This approach provides users with greater control over the audio generation process, enabling more detailed and customizable interactions with the generated audio output. Furthermore, it addresses the suboptimal performance of global V2A approaches in handling multi-object videos by transferring control to the user, thereby enhancing the overall quality and applicability of the generated audio.
To enable interactive audio generation, we develop the VGG-AnimSeg dataset, extending VGGSound with object-specific video-mask-audio triplets created via multimodal aligners (Wang et al., 2023) and video object segmentation (VOS) (Cheng et al., 2023). We then introduce Object-aware Contrastive Audio-Visual Fine-tuning (OCAV), leveraging the Mask-guided Visual Encoder (MVE) to first extract object-level visual features and then align them with audio through contrastive learning. Two data augmentation strategies, Random Video Stitching (RVS) and Mask-Guided Loudness Modulation (MLM), are developed to improve model’s sensitivity to segmented objects. Finally, we train a Latent Diffusion Model (LDM) conditioned on the features extracted by MVE to generate audio. To better assess audio-visual correspondence, we introduce a novel metric, the CAV score, for our evaluation. Our comprehensive experimental evaluations demonstrate superior performance improvements across multiple metrics.
Our main contributions are summarized as below:
@@
  • To the best of our knowledge, Hear-Your-Click is the first interactive V2A framework that allows users to generate object-specific sounds via a simple click on the videos.
@@
  • We conduct thorough experiments to validate our approach on the VGG-AnimSeg dataset which is tailored specifically for our framework. The results across multiple metrics, including the CAV score we propose, showcase state-of-the-art performance in object-level audio-visual alignment.
@@ 2.1. Video-to-Audio Generation
Significant advancements in Video-to-Audio (V2A) generation have been fueled by innovations in generative models including Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) and diffusion models (Rombach et al., 2022). Various studies propose innovative solutions to V2A challenges. SpecVQGAN (Iashin and Rahtu, 2021) trains a codebook on spectrograms to generate sounds. Im2Wav (Sheffer and Adi, 2023) employs CLIP embeddings (Radford et al., 2021) and Transformers (Vaswani, 2017) to generate audio from images. Diff-Foley (Luo et al., 2024) introduces Contrastive Audio-Visual Pre-Training (CAVP) to enhance audio-visual synchronization. Seeing and Hearing (Xing et al., 2024) introduces a multimodal latent aligner to bridge existing video and audio generation models. TiVA (Wang et al., 2024b) and MaskVAT (Pascual et al., 2024) focus on synchronization and propose the use of audio layout and sequence-to-sequence mask generative model, respectively. Other methods, such as SonicVisionLM (Xie et al., 2024), SVA (Chen et al., 2024) and FoleyCrafter (Zhang et al., 2024), use text descriptions as a mediating modality between video and audio, allowing them to generate high-quality audio using large language models (LLMs) or pre-trained text-to-audio (T2A) models, but they also face the challenge of achieving accurate synchronization (Zhang et al., 2024). The above methods focus primarily on global video information, which can lead to missing local details and difficulties in processing complex multi-object videos. And they generally lack mechanisms for fine-grained control over the V2A generation process.
@@ 3.1. Task Formulation
Given a $T$-frame video $\mathcal{V}\in\mathbb{R}^{T\times H\times W\times 3}$ and a target object $\mathcal{S}$, our objective is to generate the corresponding Mel spectrogram of the audio $\mathcal{A}$, where $\mathcal{A}\in\mathbb{R}^{T’\times N}$, with $T^{\prime}$ representing the temporal length and $N$ denoting the mel bins. To achieve this, we first obtain binary masks $\mathcal{M}$ to delineate $\mathcal{S}$, where =$\mathcal{M}=\{\mathcal{M}_{1},\mathcal{M}_{2},\ldots,\mathcal{M}_{T}\}$ and $\mathcal{M}_{t}\in\{0,1\}^{H\times W}$ refers to the $t$-th frame mask. In the inference stage, the target object $\mathcal{S}$ is specified through the user’s clicks on video frames. In the training stage, $\mathcal{S}$ is specified through textual prompts labeled by human. Leveraging the information of $\mathcal{S}$, we perform prompt-guided video segmentation to obtain $\mathcal{M}$. Then $\mathcal{M}$ together with corresponding $\mathcal{V}$ and $\mathcal{A}$ form triplets $\mathcal{(V,M,A)}$, which are used to supervise the model in learning the mappings between the input $(\mathcal{V},\mathcal{M})$ and the output $\mathcal{A}$.
Method overview. First, we tailor a new dataset named VGG-AnimSeg based on the task defined above (Sec. 3.2). Then we propose Object-aware Contrastive Audio-Visual Fine-tuning (OCAV) to align video-mask-audio triplets (Sec. 3.3), including a Mask-guided Visual Encoder (MVE) (Sec. 3.3.1) designed to extract object-level visual features, and two data augmentation strategies: Random Video Stitching (RVS) (Sec. 3.3.4) and Mask-guided Loudness Modulation (MLM) (Sec. 3.3.3). Finally, we detail the training of a latent diffusion model conditioned on the visual features obtained through OCAV (Sec. 3.4), along with the construction of the interactive inference framework (Sec. 3.5).
Figure 2. Overview of OCAV. Object-level visual features extracted by MVE are alinged with corresponding audio features. Training videos and audio tracks are augmented through (1) Random Video Stitching and (2) Mask-guided Loudness Modulation.
@@ 3.2. VGG-AnimSeg Dataset
Given that existing audio-visual datasets fail to adequately support our objective of focusing on specific objects, we have constructed a new dataset based on the VGGSound dataset (Chen et al., 2020) to address this limitation. The VGGSound dataset contains over 200K 10-second video clips spanning more than 300 categories, each accompanied by human-labeled textual descriptions. To ensure distinct vocal subjects in the videos as well as cleaner audio tracks, we initially selected all animal-related videos from VGGSound. This selection process resulted in a subset containing 68 classes of textual descriptions.
To filter out noisy samples, we use multimodal joint embeddings to select the samples where both modalities (audio and video) exhibit a strong alignment with textual descriptions. Specifically, we first leverage the Contrastive Language-Audio Pretraining (CLAP) model (Elizalde et al., 2023) and Contrastive Language-Image Pre-Training (CLIP) model (Radford et al., 2021) to extract audio, image, and text embeddings. For each sample, we compute the cosine similarity between the audio-text and image-text embedding pairs. Based on these scores, we select 400 training samples and 40 test samples per textual description with the highest average similarity, resulting in a dataset of approximately 30,000 samples.
To further ensure precise correspondence between audio and visual content, we employ DEVA (Cheng et al., 2023) to generate video binary masks via text-prompted segmentation using the text descriptions. This approach ensures a one-to-one correspondence between the audio tracks and their corresponding visual segments.
Figure 3. Enhance the correlation between object distance and volume changes.
Figure 4. Improve the model’s ability to handle multi-object scenes.
@@ 3.3. OCAV
Audio-visual alignment bridges semantic and temporal gaps between audio and visual modalities, with Contrastive Audio-Visual Pre-training (CAVP) (Luo et al., 2024) representing a notable advancement in achieving global alignment between these modalities. However, CAVP’s focus on global synchronization overlooks fine-grained, object-level details critical for precise audio-visual understanding. To address this limitation while retaining CAVP’s strengths, we propose the Object-aware Contrastive Audio-Visual Fine-tuning (OCAV) framework.
@@ 3.3.1. Mask-guided Visual Encoder (MVE)
Traditional video encoders, which process the entire video input, often fail to concentrate exclusively on the target objects indicated by the masks $\mathcal{M}$. Therefore, we propose Mask-guided Visual Encoder (MVE) to extract object-specific visual features from video-mask pairs efficiently. The MVE architecture integrates binary masks as supplementary inputs alongside the primary video stream, thereby enhancing the model’s responsiveness to the designated objects. This dual-input structure consists of two branches: a video-encoding branch $f_{v}$ and a mask-encoding branch $f_{m}$.
In the processing of the video input for $f_{v}$, we first compare feeding masked videos with original videos, and observe that masked videos lead to more stable and clearer audio as they block out background interference. This is denoted as:
@@
Regarding the mask input, we design $f_{m}$ with a convolutional backbone to capture the temporal characteristics of the masks. These extracted features are then fused with $\mathbf{\boldsymbol{x}}_{mv}$ to obtain the ultimate visual features $\mathbf{\boldsymbol{x}}_{v}$:
@@ 3.3.3. Mask-guided Loudness Modulation (MLM)
Temporal inconsistencies in the dataset, such as sounds continuing after their sources move out of the frame, create ambiguities for models. These ambiguities make it difficult for models to determine whether sounds should continue or stop once the sound sources are no longer visible. Imperfect audio-video synchronization, including off-screen sounds and background noise, complicates audio-video alignment and introduces extraneous interference to joint representation learning.
Our objective is to mitigate these issues by ensuring that the cessation of sound effects precisely corresponds to the departure of the target object from the camera view during inference. Additionally, we aim to adjust the audio loudness in accordance with the relative distances of the objects from the camera. To accomplish these objectives, we incorporate Mask-guided Loudness Modulation (MLM) into our training methodology. Through strengthening the correlation between object masks and audio loudness, MLM facilitates a more accurate learning of the mapping relationship between visual masks and auditory components.
@@ 3.3.4. Random Video Stitching (RVS)
Video augmentation methods are extensively applied in various video processing tasks, including classification, action recognition and object detection (Yun et al., 2019; Chen et al., 2023; Duan et al., 2023; Xu et al., 2024). These techniques facilitate more robust generalization and improved performance of deep learning models by generating diverse training data. Moreover, studies such as LeMDA (Liu et al., 2022) and MixGen (Hao et al., 2023) offer novel insights into multimodal data augmentation. However, such methods have been underutilized in the domain of audio-visual alignment.
In this work, we propose Random Video Stitching (RVS) during OCAV to enhance the model’s ability to identify and understand individual objects. Specifically, for each video in the dataset, another video is randomly selected, and each frame is stitched either horizontally or vertically with the corresponding frame from the original video. Simultaneously, the audio tracks from both videos are overlapped. By incorporating RVS, we aim to improve the model’s capacity to handle complex, multi-object scenes and enhance its overall performance and generalization.
Theoretically, RVS can significantly increase training samples by combining different video segments, but too many augmented samples can lead to overfitting and reduced performance on new data. Therefore, we apply RVS to the 1,126 training samples from the VGG-AnimSeg-1k dataset and integrate the augmented data into the final training set, ensuring a balance between original and augmented samples.
@@ 3.4. Latent Diffusion Model
3.5. Interative Inference
There is a growing trend in research towards interactive content generation (Geyer et al., 2023; Wang et al., 2024a; Ma et al., 2024; Li et al., 2024) because it engages users directly in the process, allowing for real-time feedback and customization. This makes the system more adaptable to specific user needs and enhances the overall user experience, providing greater enjoyment. Inspired by these studies, we develop an interactive interface for Hear-Your-Click, utilizing the promptable segmentation capabilities of the Segment Anything Model (SAM) (Kirillov et al., 2023) and the Track Anything Model (TAM) (Yang et al., 2023). This interface allows users to upload a silent video and then select a single frame to specify target regions via clicks. Upon selection, SAM generates corresponding masks in real-time based on user clicks. Users can iteratively refine these masks to ensure high precision in delineating the target object. Once the mask is finalized, TAM propagates it throughout the entire video sequence using semi-supervised video object segmentation, generating the comprehensive video mask $\mathcal{M}$. Next, we extract MVE features $\boldsymbol{x}_{v}$ and CLIP features $\boldsymbol{x}_{v}^{*}$ from $\mathcal{(V,M)}$, forming the condition embedding $\boldsymbol{x}_{c}$. Finally, we sample the final audio using trained LDM conditioned on $x_{c}$. Fig. 5 provides a detailed overview of the inference pipeline, from user interaction to audio generation.

### v2

# Hear-Your-Click: Interactive Object-Specific Video-to-Audio Generation

@@ Abstract.
Video-to-audio (V2A) generation shows great potential in fields such as film production. Despite significant advances, current V2A methods relying on global video information struggle with complex scenes and generating audio tailored to specific objects. To address these limitations, we introduce Hear-Your-Click, an interactive V2A framework enabling users to generate sounds for specific objects by clicking on the frame. To achieve this, we propose Object-aware Contrastive Audio-Visual Fine-tuning (OCAV) with a Mask-guided Visual Encoder (MVE) to obtain object-level visual features aligned with audio. Furthermore, we tailor two data augmentation strategies, Random Video Stitching (RVS) and Mask-guided Loudness Modulation (MLM), to enhance the model’s sensitivity to segmented objects. To measure audio-visual correspondence, we designed a new evaluation metric, the CAV score. Extensive experiments demonstrate that our framework offers more precise control and improves generation performance across various metrics. Project Page: https://github.com/SynapGrid/Hear-Your-Click
@@
Figure 1. Overview of the Hear-Your-Click system. After users select a single frame and interactively choose specific objects or areas, Hear-Your-Click generates the corresponding audio. This user-friendly approach empowers individuals to customize the audio according to their preferences and interests.
@@ 1. Introduction
Video-to-Audio (V2A) generation has great potential for applications in domains such as film production, social media, and accessibility services. Recent research in the field has yielded significant advances, particularly with diffusion models demonstrating remarkable performance in cross-modal generation tasks (Ruiz et al., 2023; Huang et al., 2023; Evans et al., 2024; Ma et al., 2024; Ju et al., 2024).
@@
Despite considerable advances, several challenges remain in the field of V2A generation. One key challenge is comprehending intricate semantic information within videos. Videos often contain multiple objects and actions, requiring V2A models to generate appropriate sound effects for each object and their interactions. Consequently, most V2A models struggle to handle such rich semantic information.
@@
Another challenge is the lack of fine-grained control over the audio generation process. Existing methods predominantly rely on global video information, which is insufficient for customizing audio for specific objects or regions within the video according to user needs (Iashin and Rahtu, 2021; Luo et al., 2024; Xing et al., 2024). This hinders the ability of generated audio to meet specific requirements, restricting the practical applicability of V2A models in domains like interactive media and film production. Finer control mechanisms would enhance the accuracy of generated audio and improve the feasibility of V2A models in real-world applications.
@@
To address these challenges, this work proposes an interactive V2A generation framework named Hear-Your-Click. As shown in Fig. 1, by allowing users to select specific objects within a video with a single click, Hear-Your-Click generates sounds that correspond to the selected regions, ensuring precise synchronization and alignment with the visual content. This approach provides users with greater control over audio generation, enabling more detailed and customizable interactions with the generated audio. Furthermore, it addresses the suboptimal performance of global V2A approaches with multi-object videos by transferring control to the user, enhancing the quality and applicability of the generated audio.
Our main contributions are:
@@
  • We propose Hear-Your-Click, an interactive V2A framework that allows users to generate object-specific sounds via a simple click on the videos.
@@
  • We conduct thorough experiments to validate our approach on the VGG-AnimSeg dataset which is tailored specifically for our framework. The results across multiple metrics, including our CAV score, showcase state-of-the-art performance in object-level audio-visual alignment.
@@ 2.1. Video-to-Audio Generation
Significant advancements in Video-to-Audio (V2A) generation have been fueled by innovations in generative models including Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) and diffusion models (Rombach et al., 2022). Various studies propose innovative solutions to V2A challenges. SpecVQGAN (Iashin and Rahtu, 2021) trains a codebook on spectrograms to generate sounds. Im2Wav (Sheffer and Adi, 2023) employs CLIP embeddings (Radford et al., 2021) and Transformers (Vaswani, 2017) to generate audio from images. Diff-Foley (Luo et al., 2024) introduces Contrastive Audio-Visual Pre-Training (CAVP) to enhance audio-visual synchronization. Seeing and Hearing (Xing et al., 2024) introduces a multimodal latent aligner to bridge existing video and audio generation models. TiVA (Wang et al., 2024b) and MaskVAT (Pascual et al., 2024) focus on synchronization and propose the use of audio layout and sequence-to-sequence mask generative model, respectively. Other methods, such as SonicVisionLM (Xie et al., 2024), SVA (Chen et al., 2024) and FoleyCrafter (Zhang et al., 2024), use text descriptions as a mediating modality between video and audio, allowing them to generate high-quality audio using large language models (LLMs) or pre-trained text-to-audio (T2A) models, but they also face the challenge of achieving accurate synchronization (Zhang et al., 2024). The methods above primarily focus on global video information, which can lead to missed local details and difficulties processing complex multi-object videos. Furthermore, they generally lack mechanisms for fine-grained control over the V2A generation process.
@@ 3.1. Task Formulation
Given a $T$-frame video $\mathcal{V}\in\mathbb{R}^{T\times H\times W\times 3}$ and a target object $\mathcal{S}$, our objective is to generate the corresponding Mel spectrogram of the audio $\mathcal{A}$, where $\mathcal{A}\in\mathbb{R}^{T’\times N}$, with $T^{\prime}$ representing the temporal length and $N$ denoting the mel bins. To achieve this, we first obtain binary masks $\mathcal{M}$ to delineate $\mathcal{S}$, where =$\mathcal{M}=\{\mathcal{M}_{1},\mathcal{M}_{2},\ldots,\mathcal{M}_{T}\}$ and $\mathcal{M}_{t}\in\{0,1\}^{H\times W}$ refers to the $t$-th frame mask. In the inference stage, the target object $\mathcal{S}$ is specified through user clicks on video frames. In the training stage, $\mathcal{S}$ is specified through textual prompts labeled by humans. Leveraging the information of $\mathcal{S}$, we perform prompt-guided video segmentation to obtain $\mathcal{M}$. Then $\mathcal{M}$ together with corresponding $\mathcal{V}$ and $\mathcal{A}$ form triplets $\mathcal{(V,M,A)}$, which are used to supervise the model in learning the mappings between the input $(\mathcal{V},\mathcal{M})$ and the output $\mathcal{A}$.
Method overview. First, we tailor a new dataset named VGG-AnimSeg based on the task defined above (Sec. 3.2). Then we propose OCAV to align video-mask-audio triplets (Sec. 3.3), including a Mask-guided Visual Encoder (MVE) (Sec. 3.3.1) designed to extract object-level visual features, and two data augmentation strategies: Random Video Stitching (RVS) (Sec. 3.3.4) and Mask-guided Loudness Modulation (MLM) (Sec. 3.3.3). Finally, we detail the training of a latent diffusion model conditioned on the visual features obtained through OCAV (Sec. 3.4), along with the construction of the interactive inference framework (Sec. 3.5).
Figure 2. OCAV architecture overview, showing the alignment of object-level visual features from MVE with corresponding audio features. Training data augmentation includes random video stitching and mask-guided loudness modulation.
@@ 3.2. VGG-AnimSeg Dataset
Given that existing audio-visual datasets fail to adequately support our objective of focusing on specific objects, we have constructed a new dataset based on the VGGSound dataset (Chen et al., 2020) to address this limitation. The VGGSound dataset contains over 200K 10-second video clips spanning more than 300 categories, each accompanied by human-labeled textual descriptions. To ensure distinct vocal subjects and cleaner audio, we selected all animal-related videos from VGGSound. This resulted in a subset containing 68 classes of textual descriptions.
To filter out noisy samples, we use multimodal joint embeddings to select samples where audio and video align strongly with textual descriptions. Specifically, we first leverage the Contrastive Language-Audio Pretraining (CLAP) model (Elizalde et al., 2023) and Contrastive Language-Image Pre-Training (CLIP) model (Radford et al., 2021) to extract audio, image, and text embeddings. For each sample, we compute the cosine similarity between the audio-text and image-text embedding pairs. Based on these scores, we select 400 training samples and 40 test samples per textual description with the highest average similarity, resulting in a dataset of approximately 30,000 samples.
To further ensure precise correspondence between audio and visual content, we employ DEVA (Cheng et al., 2023) to generate video binary masks via text-prompted segmentation using the text descriptions. This ensures a one-to-one correspondence between audio and corresponding visual segments.
Figure 3. Correlation enhancement between object distance and audio volume changes achieved through adaptive filtering.
Figure 4. Illustration of the improved multi-object scene handling capability of the model, achieved through a novel attention mechanism focusing on individual objects.
@@ 3.3. OCAV
Audio-visual alignment bridges semantic and temporal gaps between audio and visual modalities, with Contrastive Audio-Visual Pre-training (CAVP) (Luo et al., 2024) representing a notable advancement in achieving global alignment between these modalities. However, CAVP’s focus on global synchronization overlooks fine-grained, object-level details critical for precise audio-visual understanding. To address this limitation while retaining CAVP’s strengths, we propose the OCAV framework.
@@ 3.3.1. Mask-guided Visual Encoder (MVE)
Traditional video encoders, which process the entire video input, often fail to concentrate exclusively on the target objects indicated by the masks $\mathcal{M}$. Therefore, we propose MVE to extract object-specific visual features from video-mask pairs efficiently. The MVE architecture integrates binary masks as supplementary inputs with the primary video stream, enhancing the model’s responsiveness to designated objects. This dual-input structure consists of two branches: a video-encoding branch $f_{v}$ and a mask-encoding branch $f_{m}$.
When processing the video input for $f_{v}$, we compared feeding masked videos with original videos and observed that masked videos led to more stable and clearer audio by blocking out background interference. This is denoted as:
@@
For the mask input, we design $f_{m}$ with a convolutional backbone to capture the temporal characteristics of the masks. These extracted features are then fused with $\mathbf{\boldsymbol{x}}_{mv}$ to obtain the ultimate visual features $\mathbf{\boldsymbol{x}}_{v}$:
@@ 3.3.3. Mask-guided Loudness Modulation (MLM)
Temporal inconsistencies in the dataset, such as sounds continuing after their sources move out of the frame, create ambiguities for the models. These ambiguities make it difficult for the models to determine whether sounds should continue or stop when the sound sources are no longer visible. Imperfect audio-video synchronization, including off-screen sounds and background noise, complicates audio-video alignment and introduces interference to joint representation learning.
We aim to mitigate these issues by ensuring the cessation of sound effects precisely corresponds to the target object’s departure from the camera view during inference. Additionally, we aim to adjust audio loudness according to the objects’ relative distances from the camera. To accomplish this, we incorporate Mask-guided Loudness Modulation (MLM) into our training methodology. By strengthening the correlation between object masks and audio loudness, MLM facilitates more accurate learning of the mapping between visual masks and auditory components.
@@ 3.3.4. Random Video Stitching (RVS)
Video augmentation methods are extensively applied in various video processing tasks, including classification, action recognition and object detection (Yun et al., 2019; Chen et al., 2023; Duan et al., 2023; Xu et al., 2024). These techniques facilitate more robust generalization and improve the performance of deep learning models by generating diverse training data. Moreover, studies such as LeMDA (Liu et al., 2022) and MixGen (Hao et al., 2023) offer novel insights into multimodal data augmentation. However, such methods have been underutilized in the domain of audio-visual alignment.
We propose RVS during OCAV to enhance the model’s ability to identify and understand individual objects. Specifically, for each video in the dataset, another video is randomly selected, and each frame is stitched either horizontally or vertically with the corresponding frame from the original video. Simultaneously, the audio from both videos is overlapped. By incorporating RVS, we aim to improve the model’s capacity to handle complex, multi-object scenes and enhance its performance and generalization.
Theoretically, RVS can significantly increase training samples by combining different video segments; however, too many augmented samples can lead to overfitting and reduced performance on new data. Therefore, we apply RVS to the 1,126 training samples from the VGG-AnimSeg-1k dataset and integrate the augmented data into the final training set to ensure a balance between original and augmented samples.
@@ 3.4. Latent Diffusion Model
3.5. Interactive Inference
There is a growing trend in research towards interactive content generation (Geyer et al., 2023; Wang et al., 2024a; Ma et al., 2024; Li et al., 2024) because it engages users directly in the process, allowing for real-time feedback and customization. This makes the system more adaptable to specific user needs and enhances the user experience. Inspired by these studies, we develop an interactive interface for Hear-Your-Click, utilizing the promptable segmentation capabilities of the Segment Anything Model (SAM) (Kirillov et al., 2023) and the Track Anything Model (TAM) (Yang et al., 2023). This interface allows users to upload a silent video and then select a single frame to specify target regions via clicks. Upon selection, SAM generates corresponding masks in real time based on user clicks. Users can iteratively refine these masks to ensure precise delineation of the target object. Once the mask is finalized, TAM propagates it throughout the entire video sequence using semi-supervised video object segmentation, generating the comprehensive video mask $\mathcal{M}$. Next, we extract MVE features $\boldsymbol{x}_{v}$ and CLIP features $\boldsymbol{x}_{v}^{*}$ from $\mathcal{(V,M)}$, forming the condition embedding $\boldsymbol{x}_{c}$. Finally, we sample the audio using the trained LDM conditioned on $x_{c}$. Fig. 5 provides a detailed overview of the inference pipeline, from user interaction to audio generation.